### v2

# GAMA: Generative Agents for Multi-Agent Autoformalization

Abstract
Multi-agent simulations facilitate the exploration of interactions among both natural and artificial agents. However, modelling real-world scenarios and developing simulations often requires substantial expertise and effort. To streamline this process, we present a framework that enables the autoformalization of interaction scenarios using agents augmented by large language models (LLMs) utilising game-theoretic formalisms. The agents translate natural language descriptions of interactions into executable logic programs that define the rules of each game, ensuring syntactic correctness through validation by a solver. A tournament simulation then tests the functionality of the generated game rules and strategies. After the tournament, if a ground truth payoff matrix is available, an exact semantic validation is performed. We evaluate our approach on a diverse set of 110 natural language descriptions exemplifying five $2\times 2$ simultaneous-move games, achieving 100% syntactic and 76.5% semantic correctness in the generated game rules for Claude 3.5 Sonnet, and 99.82% syntactic and 77% semantic correctness for GPT-4o. Additionally, we demonstrate high semantic correctness in autoformalizing gameplay strategies. Overall, the results highlight the potential of autoformalization to leverage LLMs in generating formal reasoning modules for decision-making agents.
1 Introduction
Multi-agent simulations provide insights into diverse phenomena, ranging from individual human interactions and organizational dynamics to biological processes, robotics, and interactions among artificial agents. However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort. Consider, for instance, the challenge of coordinating autonomous vehicles: creating a simulation model involves specifying the rules of interaction, the goals of each agent, the information available to them, and the strategies they may employ, among other factors. Once developed, this model must then be implemented. The emergence of Large Language Models (LLMs) enables a novel method for streamlining these tasks through autoformalization. We demonstrate it using scenarios that can be modelled as $2\times 2$ simultaneous-move games.
Although LLMs have been evaluated as agents in game-theoretic contexts Akata et al. (2023); Fan et al. (2023); Guo (2023); Lorè and Heydari (2023), their inherent limitations—such as hallucinations Achiam et al. (2023) and logical or arithmetical errors Imani et al. (2023)—constrain their effectiveness as rational decision—makers in simulations. Rather than relying on LLMs as autonomous agents, a more promising approach is to harness their strengths in format translation, particularly in converting natural language into formal representations. This process, known as autoformalization Wu et al. (2022), has shown promise in formalizing mathematics Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022) and logic Yang et al. (2023); Pan et al. (2023); Cosler et al. (2023); Chen et al. (2023); Feng et al. (2023). The resulting formal representations can be utilized in combination with subsymbolic models within a neuro-symbolic framework. This approach, as demonstrated by recent advances such as AlphaGeometry, has the potential to achieve human-like problem-solving abilities in challenging domains like geometry Trinh et al. (2024).
In this work, we introduce GAMA: Generative Agents for Multi-Agent Autoformalization, a framework that leverages autoformalization to generate formal representations of strategic interactions. These representations serve as the reasoning module of agents deployed in simulations. This research builds upon our prior work Mensfelt et al. (2024), where we presented a basic framework for autoformalizing natural language interaction scenarios. In this paper, we integrate the previously developed autoformalization module into a modular agent architecture within a simulation framework. The main contributions of this work are as follows:
  1. I.
    An agent model with formal game and strategy representation: We propose an agent model that incorporates formal representations of both games and strategies, along with the autoformalization module from our earlier work. Game rules and strategies can either be autoformalized or provided using predefined rules.
  2. II.
    Three-level agent validation: We introduce a validation mechanism for agents at three levels: (1) syntactic correctness, using formal solvers; (2) functionality, by testing the code in the simulation; and (3) semantic correctness, by comparing target and achieved payoffs (when a ground truth payoff matrix is available) or verifying compliance with game type constraints (when a ground truth game type is provided).
  3. III.
    A simulation framework for autoformalized scenarios: We present a simulation framework for autoformalized interaction scenarios, enabling automated validation and facilitating the exploration and comparison of different strategies for specific scenarios.
  4. IV.
    Experimental evaluation: We evaluate the framework, demonstrating high levels of syntactic and semantic correctness for scenarios modelled as $2\times 2$ simultaneous-move games. Additionally, we assess the agents’ ability to autoformalize five distinct strategies.
2 Preliminaries
2.1 Game Theory
Game theory is an analytical framework for strategic interactions, where games are composed of four main components Rasmusen (2006):
  • Players: players consistently and willingly act towards maximising their utility, in full awareness of the situation they are involved in;
  • Actions: mappings between states and outcomes, carried out players Osborne and Rubinstein (1994);
  • Payoffs: payoffs are quantitative measures over the consequent outcomes of action tuples chosen by players Gibbons (1992);
  • Information: game information describes a player’s knowledge about the available actions, historical moves, and the player’s image of its opponents Osborne and Rubinstein (1994).
We formally define games in this context as a tuple $(N,A,\Pi)$ consisting of a set of $n$ players $N$, a non-empty set of actions $A_{i},\forall i\in N$, and for each player $i$ a payoff function $\pi_{i}:\times_{j\in N}A_{j}\rightarrow\mathbb{R}$ mapping action profiles to player payoffs.
row/col L R
U ($W_{row}$, $W_{col}$) ($X_{row}$, $X_{col}$)
D ($Y_{row}$, $Y_{col}$) ($Z_{row}$, $Z_{col}$)
Table 1: A general normal form game describing the four possible outcomes $W,X,Y,Z$, where the tuples denote the row player’s and column player’s corresponding payoffs, in general: $(\pi_{row},\pi_{col})$.
In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity. In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$. This allows us to use the terminology from Axelrod’s tournaments Axelrod and Hamilton (1981) to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
  • Prisoner’s Dilemma is characterised by a dominant temptation to defect, represented by the payoff relation $T>R>P>S$,
  • Hawk-Dove entails switched relative payoffs between the punishment for mutual defection and the sucker’s payoff – $T>R>S>P$,
  • Stag Hunt involves a relatively high reward for mutual cooperation over the temptation to defect – $R>T>P>S$.
The Battle of the Sexes and Matching Pennies are asymmetric games. As the outcomes mean different payoffs for row and column, we cannot use the generalization from above and have to define the games from both players’ perspectives:
  • Battle of the Sexes is a coordination game, where both players prefer an agreement but have different preferences over the coordinated outcomes:
    • row: $W>Z>\{X,Y\}$,
    • column: $Z>W>\{X,Y\}$.
  • Matching Pennies: coordinated outcomes reward the row player, while mismatched actions benefit the column player:
    • row: $\{W,Z\}>\{X,Y\}$,
    • column: $\{X,Y\}>\{W,Z\}$.
2.2 Large Language Models
The rapid advancement of natural language processing (NLP), driven by transformer architectures Gillioz et al. (2020) and pre-trained models Qiu et al. (2020), has led to the emergence of Large Language Models (LLMs). State-of-the-art (SOTA) LLMs continue to evolve rapidly, primarily due to the increasing scale of model parameters and training datasets Zhao et al. (2023).
In this work, we employ OpenAI’s GPT-4 Omni OpenAI (2024), a multimodal SOTA model capable of processing text, images, and audio. It has demonstrated strong performance on reasoning benchmarks Wang and Zhao (2024). Likewise, we utilize Anthropic’s Claude 3.5 Sonnet, another high-performing SOTA model, which excels across multiple benchmarks Anthropic (2024).
2.3 General Game Playing
General game playing Genesereth et al. (2005) focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention. The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games Love et al. (2006). GDL focused on information games only, so it was extended in GDL-II Thielscher (2010) to cover n-player games with incomplete information and games in extensive normal form Thielscher (2011).
A challenge with GDL systems is that learning without human guidance demands complex reasoning. Players must infer the possible actions of others, effectively analysing hypothetical game scenarios before making decisions. Action formalisms like the classical Situation Calculus McCarthy and Hayes (1981) have been developed for precisely this purpose. Formal schemes and inference methods are readily available for Situation Calculus Giacomo et al. (2010); Lesperance et al. (2024), while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages. One such scheme Schiffel and Thielscher (2011) shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents Scherl and Levesque (2003). Our game solver is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
3 Generative Agents for Multi-Agent Autoformalization
3.1 Overview
Figure 1: GAMA framework: each agent’s game rules and strategy can be predefined, autoformalized, or loaded from an existing agent. The initial agent pool undergoes syntactic validation, and agents that fail to produce syntactically valid code within a defined number of attempts are eliminated (denoted by a crossed circle). The remaining agents are validated in the tournament and subsequently undergo semantic validation.
The GAMA framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents (Fig. 1). An agent consists of an autoformalization module and a logic programming component implemented in Prolog, which is divided into two modules: a Game module, containing both game-independent and game-dependent rules, and a Strategy module, which describes the actions to select against an opponent. Together, these modules form a solver that encapsulates the game logic and determines the agent’s next move. A Python wrapper facilitates the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state. This modular design allows for easy substitution of game-dependent rules and strategies, either through predefined configurations or by autoformalizing natural language descriptions. The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/GAMA.
3.2 Solver
The solver comprises a game-independent part that represents the rules for any extensive-form game, a game-dependent part that defines the rules of a particular game using the predicates from the game-independent part, and a strategy part to select an action. Our solver is implemented in Prolog Körner et al. (2022); Wielemaker et al. (2012)—a declarative programming language well-suited for high-level problem definitions—but it can easily be substituted with another formal solver. The game-independent and game-dependent parts were described in detail in Mensfelt et al. (2024). To enable the agent incorporating the solver to play, we extended it with the strategy part. Additionally, we implemented semantic validation in Prolog using validity constraints. The strategy part and constraint-checking mechanism are described below.
3.2.1 Strategy
Strategies are submitted separately from the game-dependent component. The select/4 predicate is used to determine the next move M of player P against opponent O in game situation S, as illustrated in the tit-for-tat strategy example below:
select(P, O, S, M) if
not holds(last_move(O, _LMo), S) and
holds(default_move(P, M), S).
select(P, O, S, Mo) if
holds(last_move(O, Mo), S).
Here, a player P selects the default move first (typically to cooperate), otherwise it mirrors the opponent’s move in the last round. holds/2 is based on the implementation of the Situation Calculus McCarthy and Hayes (1981). The strategy relies on specific predicates – default_move/2 and last_move/2 – which are required from the game-dependent part of the solver. More sophisticated strategies can be expressed.
3.2.2 Validity checking
We check that the auto-formalized payoff matrix satisfies the structural and content constraints of the specified game (as outlined in Sect. 2.1), by defining validity constraints. The listing below specifies the payoff matrix validity constraints for the Prisoner’s Dilemma.
valid_pd_payoffs(T,R,P,S,C,D) if
payoff(C,C,R,R) and
payoff(C,D,S,T) and
T>R and
payoff(D,C,T,S) and
payoff(D,D,P,P) and
R>P and
P>S.
Since the autoformalization module does not impose constraints on the names of actions, this predicate also identifies which actions names correspond to specific actions.
3.3 Autoformalization Module
Algorithm 1 Generating game rules predicates from natural language descriptions. PD stands for Prisoner’s Dilemma.
1:Input:
$\Gamma$: game-independent predicates,
$NL_{PD}$: natural language description of PD,
$\xi_{PD}$: game-specific predicates of PD,
$NL_{NG}$: natural language description of a new game.
2:Output:
$\xi_{NG}$: game-specific predicates for the new game.
3:Parameter:
max_attempts: maximum correction attempts.
4:$\text{attempts}\leftarrow 0$
5:$\text{trace}\leftarrow\emptyset$
[…]
7:    $\xi_{NG}\leftarrow\text{LLM.translate}(\Gamma,NL_{PD},\xi_{PD},NL_{NG})$
[…]
9:    if is_valid then
10:         return $\xi_{NG}$
11:    else
[…]
13:         $\xi_{NG}\leftarrow\text{LLM.self\_correct}(\xi_{NG},\text{trace})$
14:    end if
15:    $\text{attempts}\leftarrow\text{attempts}+1$
16:end while
17:return Unable to generate valid predicates within maximum attempts.
To perform autoformalization, a large language model is employed. The framework supports the use of any language model by providing an abstract interface. Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma or the tit-for-tat strategy as examples, respectively. After generating the code, its syntactic correctness is validated. If errors are detected, the LLM receives feedback, including the erroneous lines, error messages, and instructions for correcting the predicates. The max_attempts parameter specifies the maximum number of correction attempts. The overview of autoformalization algorithm is presented in Listing 1 Mensfelt et al. (2024).
3.4 Simulation
The framework is employed in this work in two primary modes: validation and simulation. These modes facilitate the evaluation of both the semantic correctness of generated programs and the effectiveness of strategies within various game scenarios.
In validation mode, the framework ensures the semantic correctness of generated programs by comparing their performance against predefined criteria. Once a scenario is autoformalized, an agent is created and provided with a specified strategy. The agent then plays against a clone of itself for a set number of rounds. If a ground truth payoff matrix is available, the framework performs semantic validation by verifying that the payoff for each unique combination of moves matches the target payoff for that combination. Additionally, if the ground truth type of the game is known, the framework enforces constraint checking to confirm that the payoff matrix adheres to all specified constraints for the given game.
In simulation mode, the framework evaluates and compares strategies to determine their effectiveness within a game scenario. Agents equipped with different strategies compete in a round-robin tournament, where each agent plays against every other. By simulating multiple rounds, the framework identifies the strategy that performs best on average, offering insights into the relative strengths and weaknesses of competing approaches.
The parameters of the tournament are as follows:
  • Agent Pool: Contains all agents participating in the tournament. Each agent can be equipped with game rules and strategies either through autoformalization or predefined configurations.
  • Number of Rounds: Specifies how many rounds each agent will play.
  • Match Maker: A function that determines the pairing of agents for matches.
  • Target Payoffs: Optional parameter used for semantic validation.
The framework provides two methods for determining the tournament’s outcome:
  1. 1.
    Selecting agents whose total payoff matches their target payoff, applicable when target payoffs are specified (validation).
  2. 2.
    Selecting agents that achieve the highest total payoff, suitable for assessing the success of strategies (simulation).
Furthermore, the customizable Match Maker function allows for experiments with tailored matching protocols
4 Methods
4.1 Parameters
Parameter Value
LLM Claude 3.5 Sonnet
GPT-4o
temperature $1$
maximum output tokens $2048$
maximum attempts number $5$
Table 2: Common experimental parameters.
Parameter Exp. 1 Exp. 2 Exp. 3 Exp. 4
games $55$ $55$ $5$ $5$
agents in game $5$ $5$ $1$ $1$
strategies $1$ $1$ $6$ $6$
rounds $4$ $4$ $10$ $10$
match maker S S RR S
Table 3: Experiment-specific parameters. S stands for a mode in which an agent plays against its clone with an opposite strategy, while RR stands for round robin.
To evaluate the framework, we conducted four experiments: i) autoformalization of game scenarios with numerical (Exp. 1) and ii) non-numerical (Exp. 2) payoffs, iii) using the autoformalized game rules in a round-robin tournament (Exp.3), and iv) autoformalization of strategies (Exp. 4). Table 2 provides the experimental parameters shared across all experiments, while Table 3 details the parameters unique to each specific experiment.
4.2 Autoformalization of Interaction Descriptions (Numerical Payoffs)
We used a dataset of 55 natural language game-theoretic scenarios with numerical payoffs to evaluate the syntactic and semantic accuracy of generated logic programs. The dataset included 5 scenarios that employed common metaphors (e.g., two prisoners in the Prisoner’s Dilemma) for the five canonical games introduced in Section 2.1. The remaining 50 scenarios—10 for each game—used alternative descriptions that avoided these typical metaphors (e.g., two politicians competing in a campaign rather than two prisoners in the Prisoner’s Dilemma). For each scenario, five agents autoformalized game rules and then played tournament against their clones, generating all four possible pairs of moves in four rounds. Each agent had a maximum of five attempts to produce syntactically correct code; agents failing after five attempts were labelled as syntactically incorrect. Post-tournament, each agent’s individual and total payoffs were compared to the target, with matches indicating semantic correctness.
4.3 Autoformalization of Interaction Descriptions (Non-numerical Payoffs)
This experiment was analogous to the previous one, but it utilized interaction scenarios without numerical payoffs (e.g., ”If both remain silent, they get minor sentences,” to describe the outcome of mutual cooperation in PD). Therefore, the autoformalizing LLM had to infer the appropriate payoffs correctly. Since it was not possible to provide a ground truth matrix in this case, we validated the semantic correctness by performing constraint checks (described in Sect. 3.2) to determine whether the payoff matrix satisfied the constraints for the ground truth game type (Sect. 2.1).
4.4 Axelrod’s Tournament
Strategy Description
anti-default-move Always select the move that is the opposite of the default move.
anti-tit-for-tat Start with a default move. Then, select the move that is the opposite of the opponent’s move in the previous round.
best-response Start with a default move. Then, select a move that would give you the highest payoff in response to the opponent’s move in the previous round.
default-move Always select the default move.
random Select one of the possible moves with uniform probability.
tit-for-tat Start with a default move. Then, mirror the opponent’s move in the previous round.
Table 4: Strategies utilized in the evaluation.
To demonstrate the framework’s usefulness in analysing interaction scenarios, we selected five agents created during the first experiment, one for each of the five games. For each game, we conducted a tournament modelled after Axelrod’s tournament Axelrod and Hamilton (1981). In this setup, six copies of each agent were created and assigned one of six strategies. Each agent then played against every other agent, including itself. The strategies used are listed in Table 4.
The main metric for evaluating strategy performance was the total payoff achieved by an agent using a given strategy in each game. To allow for cross-game comparisons of strategy effectiveness, total payoffs were normalized using:
$norm\_payoff=\frac{total\_payoff-min\_total\_payoff}{{max\_total\_payoff-min\_total\_payoff}},$
$Total\_payoff$ refers to the sum of the payoffs an agent collects across all rounds and $min\_total\_payoff$ and $max\_total\_payoff$ represent the minimum and maximum total payoffs achievable, respectively.
4.5 Autoformalization of Strategies
We then evaluated the semantic correctness of the autoformalized strategies. To achieve this, we used the natural language description of the tit-for-tat strategy, along with its Prolog implementation, as an example. For the other five strategies, the autoformalization module was provided with their natural language descriptions, similar to those shown in Table 4.
Each autoformalized strategy was then assigned to an agent with predefined Prisoner’s Dilemma game rules. This agent played four rounds against a clone utilizing the anti-tit-for-tat strategy. The agent’s total payoff was compared to a target payoff specific to each strategy to assess correctness. For the random strategy, correctness was also manually validated. This tournament setup was repeated five times.
5 Results and Discussion
5.1 Autoformalization of Interaction Descriptions
1 2 3 4 5
Numeric payoffs
Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
GPT-4o 96.77% 2.87% 0.00% 0.00% 0.36%
Non-numeric payoffs
Claude 3.5 S 100.00% 0.00% 0.00% 0.00% 0.00%
GPT-4o 97.64% 2.02% 0.34% 0.00% 0.00%
Table 5: Distribution of attempts at creating syntactically correct code for each model.
Table 5 shows the distribution of attempts needed to generate syntactically correct code for both models. The type of description (numeric vs. non-numeric) had no influence on syntactic correctness of the generated code. Claude 3.5 Sonnet achieved syntactically correct code in 100% of cases and GPT-4o in nearly 100%. The few errors observed involved non-Prolog comment characters and non-escaped special characters in strings, all of which were successfully corrected on the second attempt. This demonstrates that providing a specific line that generated the error in the feedback prompt significantly enhances an LLM’s ability to self-correct compared to merely providing an error message Mensfelt et al. (2024). The only case where GPT-4o was unable to correct a syntactic error involved the error message “Unexpected end of file“ which indicated the end of the file as the error’s source. However, the actual line generating the error was not provided in the feedback prompt.
Syntactic Runtime Semantic
Numeric payoffs
Claude 3.5 Sonnet 100.0% 88.0% 86.0%
GPT-4o 100.0% 90.0% 86.0%
Non-numeric payoffs
Claude 3.5 Sonnet 100.0% 79.0% 67.0%
GPT-4o 100.0% 91.0% 68.0%
Table 6: Syntactic, runtime, and semantic correctness for both models and types of scenario descriptions.
The type of description had no influence on runtime errors in GPT-4o (runtime correctness of 90% and 91% for numeric and non-numeric payoffs, respectively, Table 6). However, it did have an influence in the case of Claude 3.5 Sonnet, with a 9 percentage point decrease to 79% for non-numerical payoffs. A runtime error in this context means that the program was syntactically correct; however, while querying a predicate during runtime—most often to calculate payoffs—a predicate necessary to execute the query was missing. This highlights the need to extend the autoformalization feedback loop to the runtime stage and leverage the tracing mechanism of the solver to identify and create missing predicates.
Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 89.1% 96.4% 68.5% 96.4% 87.3%
Semantic 89.1% 96.4% 68.5% 89.1% 87.3%
Non-Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 82.1% 85.7% 74.0% 70.7% 83.1%
Semantic 82.1% 77.8% 34.2% 64.0% 83.1%
Table 7: Correctness by game for Claude 3.5 Sonnet for game descriptions with numeric and non-numeric payoffs.
Numeric Payoffs
BS HD MP PD SH
Syntactic 98.3% 100.0% 100.0% 100.0% 100.0%
Runtime 83.1% 89.1% 96.4% 89.1% 90.9%
Semantic 79.7% 89.1% 90.9% 85.5% 83.6%
Non-Numeric Payoffs
BS HD MP PD SH
Syntactic 100.0% 100.0% 100.0% 100.0% 100.0%
Runtime 93.2% 88.1% 87.3% 96.5% 91.5%
Semantic 93.2% 67.8% 30.2% 73.7% 76.3%
Table 8: Correctness by game for GPT-4o for game descriptions with numeric and non-numeric payoffs.
The semantic correctness was lower than runtime correctness for both models and description types (Table 6). This means that even though an agent was able to select a move during runtime and update its state, the semantics of the program were not correct. For descriptions with numerical payoffs, this implies that the payoff matrix does not contain the required payoffs. For descriptions without numerical payoffs, it indicates that the payoff matrix does not satisfy the constraints of the required game. Tables 7 and 8 present the syntactic, runtime, and semantic correctness for different games.
The runtime and semantic correctness varied depending on the model and the game. Notably, Matching Pennies consistently resulted in low semantic correctness for non-numerical descriptions, with 34.2% correctness for Claude 3.5 Sonnet and 30.2% correctness for GPT-4o. This may be attributed to some descriptions specifying which player wins if the choices match or mismatch, without explicitly stating that the other player will lose—a constraint of the game. Extending the autoformalization feedback loop to include semantic evaluation has the potential to significantly improve correctness.
5.2 Axelrod’s Tournament
Figure 2: Normalized total payoff for each strategy by game.
Figure 2 presents the normalized total payoffs for each strategy across all games. Overall, the best-response strategy proved to be the most effective on average, followed closely by tit-for-tat. Best-response was most effective in all games except the Battle of the Sexes and the Prisoner’s Dilemma. In the Prisoner’s Dilemma, best-response achieved a lower payoff than the anti-default-move strategy (analogous to “defect”), as it initiated with the “cooperate” move. The results demonstrate that agents with autoformalized rules can effectively simulate and assess the performance of various strategies within different modelled scenarios.
5.3 Autoformalization of Strategies
Strategy Claude 3.5 Sonnet GPT-4o
anti-default-move 100.00% 100.00%
anti-tit-for-tat 100.00% 100.00%
best-response 80.00% 100.00%
default-move 100.00% 100.00%
random 80.00% 100.00%
Table 9: Percentage of semantic correctness for each strategy.
Table 9 presents the percentage of correctly autoformalized strategies. GPT-4o achieved 100% correctness in autoformalizing all five strategies. Claude 3.5 Sonnet produced one incorrect strategy for best-response and one for random, which are the strategies that deviate the most from the tit-for-tat strategy presented as an example in the prompt. One error occurred during runtime, and another during semantic validation. Overall, the high correctness demonstrates that the GAMA framework has the potential to effectively autoformalize both game rules and gameplay strategies.
6 Related Work
Originally proposed for agents in game scenarios, the Game Description Language (GDL) offers a promising approach for formalising complex strategic interactions in multi-agent environments by providing a declarative framework to define game rules, legal moves, and outcome conditions Genesereth et al. (2005). Agents that employ GDL are often used to systematically explore game states and strategic options, supporting rigorous analysis across both cooperative and competitive games, as suggested by previous approaches (e.g., see  Schiffel and Thielscher (2009)). However, these approaches typically rely on human-driven formalisation of specific games, a limitation addressed by the LLM-augmented agents and the GDL representation used here, which enable autoformalization and enhance flexibility.
There is a growing body of research exploring the use of LLMs in game theory Lorè and Heydari (2024), systematically examining their potential as player agents in game-theoretic scenarios Fan et al. (2024), utilising their inference abilities to make decisions in games with imperfect information Guo et al. (2023), comparing their behaviour with human participants Fontana et al. (2024), and deploying them in simulation frameworks Mao et al. (2024). While we share with these works an interest in the broad capabilities and value of LLMs in game theoretic settings, our focus diverges by concentrating on how to use LLMs to generate GDL-like executable game specifications. These formal specifications enable agents to rigorously analyse game situations and reason about possible outcomes, providing a structured and robust foundation for decision-making within game-theoretic frameworks. Our work builds on previous research discussed in Mensfelt et al. (2024), significantly extending it by embedding the autoformalization module within an agent model, incorporating the autoformalization of player strategies, implementing a distinct correctness validation process for the autoformalization through tournament play, and conducting experiments to validate the results.
Autoformalization Wu et al. (2022), has shown promising results in translating natural language into mathematical formalisms, enabling the integration of large language models with formal tools for solving reasoning tasks Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022). This approach has also been successfully extended to creating logical representations from natural language descriptions. For instance, Yang et al. (2023) applied GPT-3 with few-shot prompting to translate natural language sentences into formal representations, which were then inputted into answer set programs. While effective on NLP benchmarks, this method faced occasional translation errors. To further enhance performance, Pan et al. (2023) combined in-context learning with self-refinement based on solver evaluations, surpassing standard LLM and chain-of-thought prompting approaches on logical reasoning benchmarks. In the context of temporal logic, Cosler et al. (2023) addressed the problem of translation errors with an interactive system allowing users to refine sub-translations. Fine-tuning LLMs has also proven effective for boosting accuracy in translations to temporal logics Chen et al. (2023) and deductive datasets Feng et al. (2023).
7 Conclusions and Future Work
We have presented a framework for the autoformalization of interaction scenarios, specifically those modelled by $2\times 2$ simultaneous-move games. The framework enables the autoformalization of strategies for these games and the simulation of tournaments based on the autoformalized rules. The generated programs undergo three levels of validation: syntactic validation by consulting a solver, functional testing through gameplay, and semantic validation by comparing achieved payoffs with target outcomes or by checking the game constraints. Evaluation of the framework demonstrated high syntactic and semantic accuracy in the code generated by agents across various interaction scenarios.
Future work will focus on extending the framework to model games beyond $2\times 2$ simultaneous-move scenarios. Our previous work Mensfelt et al. (2024) demonstrated that using our sample solver for the Prisoner’s Dilemma we could successfully generate code for structurally similar games, such as rock-paper-scissors (with a different number of moves) and sequential versions of the PD. However, achieving successful generalization to a broader class of games will require assembling a more diverse set of in-context learning examples and developing methods to retrieve the most relevant examples for autoformalization in specific scenarios. Another direction for future work involves extending the autoformalization feedback loop to include runtime and semantic validation checks, further improving the correctness of the generated programs.
These extensions will provide a robust framework for generating reasoning modules for decision-making agents in a wider class of games and game-like interactions.
Acknowledgments
This work was supported by a Leverhulme Trust International Professorship Grant (LIP-2022-001). The first author would like to thank Bartosz Kosowski for the insightful discussions that contributed to the development of this research.